# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ 5.3.2 Step 2: Analysis of $W_{n+1}$
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}$ $=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(U^{i,j})-\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})|^{2}$
$=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(X_{n}^{i,\tau_{i}(j)})-\bar{V}_{i}^{\prime}(X_{n}^{i,\tau_{i}(j)},q_{\bm{X}_{n}}^{-i})|^{2}$
$=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(X_{n}^{i,j})-\bar{V}_{i}^{\prime}(X_{n}^{i,j},q_{\bm{X}_{n}}^{-i})|^{2}.$
@@
$\mathbb{E}\Big{(}|g_{1}(X_{n}^{1,j})-\bar{V}_{1}^{\prime}(X_{n}^{1,j},q_{\bm{X}_{n}}^{-1})|^{2}\Big{|}\bm{X}_{n}\Big{)}$ $=\operatorname{{\rm var}}[g_{1}(X_{n}^{1,j})|\bm{X}_{n}]=\frac{1}{B}\operatorname{{\rm var}}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})|\bm{X}_{n}]$
$\leq\frac{1}{B}\mathbb{E}\Big{\{}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})]^{2}\Big{|}\bm{X}_{n}\Big{\}}.$
@@
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}|M^{1,j}-W^{1,j}|^{2}\leq\frac{h^{2}}{BN}\sum_{j=1}^{N}\mathbb{E}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})]^{2}=\frac{h^{2}}{B}\mathbb{E}|\partial_{1}V(\bm{z}_{n}^{:,1})|^{2}.$
@@
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}\leq\frac{h^{2}}{B}\mathbb{E}\|\nabla V(\bm{z}_{n}^{:,1})\|_{2}^{2}=\frac{h^{2}}{B}\mathbb{E}_{\bm{X}_{n}}[\mathbb{E}_{\bm{z}\sim q_{\bm{X}_{n}}}\|\nabla V(\bm{z})\|_{2}^{2}].$ (5.4)
We now relate the right-hand side of (5.4) to $\mathbb{E}_{\bm{y}\sim q_{\bm{Y}}}\|\nabla V(\bm{y})\|_{2}^{2}$. We condition on $(\bm{X}_{n},\bm{Y})$ and denote by $\gamma$ an $\mathcal{W}_{2}$-optimal coupling between $q_{\bm{X}_{n}}$ and $q_{\bm{Y}}$ such that
$\mathbb{E}_{(\bm{u},\bm{v})\sim\gamma}\|\bm{u}-\bm{v}\|_{2}^{2}=\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}}).$
Construct random vectors $\bm{z},\bm{y}\in\mathbb{R}^{m}$ such that the conditional distribution of $(\bm{z},\bm{y})$ given $(\bm{X}_{n},\bm{Y})$ is $\gamma$. Then,
$\mathbb{E}\|\bm{z}-\bm{y}\|_{2}^{2}=\mathbb{E}[\mathbb{E}(\|\bm{z}-\bm{y}\|_{2}^{2}|\bm{X}_{n},\bm{Y})]=\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}}).$
@@
$\leq 2L^{2}\mathbb{E}\|\bm{z}-\bm{y}\|_{2}^{2}+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}$
$\leq 2L^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}.$
@@
The conditional distribution of $\bm{y}$ given $\bm{Y}$ is $q_{\bm{Y}}$. Hence, $\bm{y}\sim q_{*}$ and $\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}=\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}\leq mL^{2}/\alpha$. Plugging these estimates into (5.4), we get
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|M^{i,j}-W^{i,j}|^{2}\leq\frac{2h^{2}L^{2}}{B}[\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+m/\alpha].$ (5.5)
5.3.4 Step 4: Analysis of $\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}$
@@
$Y_{h}^{i,j}$ $=Y_{0}^{i,j}-\int_{0}^{h}\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})\mathrm{d}t+\sqrt{2}B_{h}^{i,j}$
$=[Y^{i,j}-h\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i})]-\int_{0}^{h}[\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})]\mathrm{d}t+\sqrt{2}B_{h}^{i,j}.$
Then,
$W^{i,j}-Y_{h}^{i,j}$ $=\Big{(}\underbrace{[U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})]}_{Y_{+}^{i,j}}-\underbrace{[Y^{i,j}-h\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i})]}_{U_{+}^{i,j}}\Big{)}$
$\quad-\underbrace{\int_{0}^{h}[\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})]\mathrm{d}t}_{A^{i,j}}.$
We have
$\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}=\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}.$ (5.6)
We make two claims and defer their proofs to Sections 5.4 and 5.5, respectively.
Claim 5.1.
For any $\varepsilon>0$,
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}]+L^{2}h^{3}.$
Claim 5.2.
For any $j\in[N]$,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq(1-\alpha h)^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})$
By (5.6) and Claim 5.1,
$\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}\leq(1+\varepsilon h)\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}]+L^{2}h^{3}.$
Based on Claim 5.2,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}$
$\leq(1+\varepsilon h)\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}+m\widetilde{L}^{2}]+mL^{2}h^{3}$
$\leq(1+\varepsilon h)(1-\alpha h)^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+h^{3}m[(\varepsilon^{-1}+h)(L^{4}/\alpha+\widetilde{L}^{2})+L^{2}].$
Take $\varepsilon=\alpha$. We have, $\varepsilon^{-1}+h=\alpha^{-1}+2/(\alpha+L)\leq 2\alpha^{-1}$ and thus,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|W^{i,j}-Y_{h}^{i,j}|^{2}\leq(1-\alpha h)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+mh^{3}[2\alpha^{-1}(L^{4}/\alpha+\widetilde{L}^{2})+L^{2}].$ (5.7)
5.3.5 Step 5: Establishing the recursive inequality
By (5.3), (5.5), (5.7) and $h\leq B\alpha/(4L^{2})$,
$W_{n+1}^{2}\leq\bigg{(}1-\alpha h+\frac{2L^{2}h^{2}}{B}\bigg{)}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+mh^{3}[2\alpha^{-1}(L^{4}/\alpha+\widetilde{L}^{2})+L^{2}]+\frac{2mh^{2}L^{2}/\alpha}{B}$
$\leq\bigg{[}1-\alpha h\bigg{(}1-\frac{2L^{2}h}{B\alpha}\bigg{)}\bigg{]}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+mh^{2}\max\{h,(LB)^{-1}\}\bigg{(}\frac{2(L^{4}/\alpha+\widetilde{L}^{2})}{\alpha}+L^{2}+L^{3}/\alpha\bigg{)}$
$\leq(1-\alpha h/2)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})+4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha).$
By (5.1),
$(1-\alpha h/2)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}})$ $\leq(1-\alpha h/4)^{2}\bigg{(}W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}$
$\leq\bigg{(}(1-\alpha h/4)W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}.$
Therefore,
$W_{n+1}^{2}\leq\bigg{(}(1-\alpha h/4)W_{n}+2\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})}\bigg{)}^{2}+4mh^{2}\max\{h,(LB)^{-1}\}(L^{4}/\alpha^{2}+\widetilde{L}^{2}/\alpha).$
The proof is completed by combining the above estimates.

### v3

@@ 5.5.2 Step 2: Error decomposition
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}=\frac{h^{2}}{N}\sum_{j=1}^{N}\mathbb{E}|g_{i}(X_{n}^{i,j})-\bar{V}_{i}^{\prime}(X_{n}^{i,j},q_{\bm{X}_{n}}^{-i})|^{2}.$
@@
$\mathbb{E}\Big(|g_{1}(X_{n}^{1,j})-\bar{V}_{1}^{\prime}(X_{n}^{1,j},q_{\bm{X}_{n}}^{-1})|^{2}\Big|\bm{X}_{n}\Big)=\operatorname{{\rm var}}[g_{1}(X_{n}^{1,j})|\bm{X}_{n}]$
$=\frac{1}{B}\operatorname{{\rm var}}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})|\bm{X}_{n}]\leq\frac{1}{B}\mathbb{E}\Big\{[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})]^{2}\Big|\bm{X}_{n}\Big\}.$
@@
$\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}|X_{n+1}^{1,j}-W^{1,j}|^{2}\leq\frac{h^{2}}{BN}\sum_{j=1}^{N}\mathbb{E}[\partial_{1}V(X_{n}^{1,j},\bm{z}_{n}^{-1,1})]^{2}=\frac{h^{2}}{B}\mathbb{E}|\partial_{1}V(\bm{z}_{n}^{:,1})|^{2}.$
@@
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|X_{n+1}^{i,j}-W^{i,j}|^{2}\leq\frac{h^{2}}{B}\mathbb{E}\|\nabla V(\bm{z}_{n}^{:,1})\|_{2}^{2}=\frac{h^{2}}{B}\mathbb{E}_{\bm{X}_{n}}[\mathbb{E}_{\bm{z}\sim q_{\bm{X}_{n}}}\|\nabla V(\bm{z})\|_{2}^{2}].$ (5.9)
We now relate the right-hand side of (5.9) to $\mathbb{E}_{\bm{y}\sim q_{\bm{Y}}}\|\nabla V(\bm{y})\|_{2}^{2}$. We condition on $(\bm{X}_{n},\bm{Y}_{nh})$ and denote by $\gamma$ an $\mathcal{W}_{2}$-optimal coupling between $q_{\bm{X}_{n}}$ and $q_{\bm{Y}_{nh}}$ such that
$\mathbb{E}_{(\bm{u},\bm{v})\sim\gamma}\|\bm{u}-\bm{v}\|_{2}^{2}=\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}}).$
Construct random vectors $\bm{z},\bm{y}\in\mathbb{R}^{m}$ such that the conditional distribution of $(\bm{z},\bm{y})$ given $(\bm{X}_{n},\bm{Y}_{nh})$ is $\gamma$. Then,
$\mathbb{E}\|\bm{z}-\bm{y}\|_{2}^{2}=\mathbb{E}[\mathbb{E}(\|\bm{z}-\bm{y}\|_{2}^{2}|\bm{X}_{n},\bm{Y}_{nh})]=\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}}).$
@@
$\leq 2L^{2}\mathbb{E}\|\bm{z}-\bm{y}\|_{2}^{2}+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}\leq 2L^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})+2\mathbb{E}\|\nabla V(\bm{y})\|_{2}^{2}.$
@@
The conditional distribution of $\bm{y}$ given $\bm{Y}_{nh}$ is $q_{\bm{Y}_{nh}}$. Hence, $\bm{y}\sim q_{*}$. Plugging the above estimate into (5.9), we get (5.8).
5.5.4 Step 4: Analysis of $\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}$
We will show that
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}\leq$ $(1-\alpha h)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})$
$+\frac{4h^{2}L^{2}}{\alpha}[h\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}+m].$ (5.10)
The error bound consists of a contraction term and an $O(h^{2})$ additive term.
To prove the result, we first use (5.4) to get
$Y_{(n+1)h}^{i,\tau_{n,i}(j)}=Y_{nh}^{i,\tau_{n,i}(j)}-\int_{nh}^{(n+1)h}\bar{V}_{i}^{\prime}(Y_{t}^{i,\tau_{n,i}(j)},q_{*}^{-i})\mathrm{d}t+\sqrt{2}(B_{(n+1)h}^{i,j}-B_{nh}^{i,j})$
$=[Y_{nh}^{i,\tau_{n,i}(j)}-h\bar{V}_{i}^{\prime}(Y_{nh}^{i,\tau_{n,i}(j)},q_{*}^{-i})]$
$\quad-\int_{nh}^{(n+1)h}[\bar{V}_{i}^{\prime}(Y_{t}^{i,\tau_{n,i}(j)},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{nh}^{i,\tau_{n,i}(j)},q_{*}^{-i})]\mathrm{d}t+\sqrt{2}(B_{(n+1)h}^{i,j}-B_{nh}^{i,j}).$
Then,
$W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}$ $=\Big(\underbrace{[X_{n}^{i,j}-h\bar{V}_{i}^{\prime}(X_{n}^{i,j},q_{\bm{X}_{n}}^{-i})]}_{U_{+}^{i,j}}-\underbrace{[Y_{nh}^{i,\tau_{n,i}(j)}-h\bar{V}_{i}^{\prime}(Y_{nh}^{i,\tau_{n,i}(j)},q_{*}^{-i})]}_{Y_{+}^{i,j}}\Big)$
$\quad-\underbrace{\int_{nh}^{(n+1)h}[\bar{V}_{i}^{\prime}(Y_{t}^{i,\tau_{n,i}(j)},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{nh}^{i,\tau_{n,i}(j)},q_{*}^{-i})]\mathrm{d}t}_{A^{i,j}}.$
We have
$\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}=\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}.$ (5.11)
We make two claims and defer their proofs to Sections˜5.6 and 5.7, respectively.
Claim 5.1.
Choose any $\varepsilon>0$. Under Assumption 4.1, we have
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+2(\varepsilon^{-1}+h)L^{2}h^{2}[h\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+1].$
Claim 5.2.
Let Assumption 4.1 hold and $0<h<2/(\alpha+L)$. For any $j\in[N]$,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq(1-\alpha h)^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})$
By (5.11) and Claim 5.1,
$\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}\leq(1+\varepsilon h)\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+2(\varepsilon^{-1}+h)L^{2}h^{2}[h\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+1].$
Based on Claim 5.2,
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|W^{i,j}-Y_{(n+1)h}^{i,\tau_{n,i}(j)}|^{2}$
$\leq$ $\frac{1+\varepsilon h}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+2(\varepsilon^{-1}+h)L^{2}h^{2}[h\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}+m]$
$\leq$ $(1+\varepsilon h)(1-\alpha h)^{2}\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})+2(\varepsilon^{-1}+h)L^{2}h^{2}[h\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}+m].$
Next, take $\varepsilon=\alpha$. Since $h<2/(\alpha+L)\leq 1/\alpha$, we have $\varepsilon^{-1}+h\leq 2/\alpha$. Then, the result (5.10) becomes obvious.
5.5.5 Step 5: Proving the recursive inequality
By (5.7), (5.8) and (5.10),
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n+1}},q_{\bm{Y}_{(n+1)h}})$ $\leq\bigg(1-\alpha h+\frac{2L^{2}h^{2}}{B}\bigg)\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{\bm{Y}_{nh}})$
$\quad+\bigg(\frac{2h^{2}}{B}+\frac{4h^{3}L^{2}}{\alpha}\bigg)\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}+\frac{4mh^{2}L^{2}}{\alpha}.$
The assumption $h<\frac{B\alpha}{4L^{2}}$ implies $2L^{2}h^{2}/B<\alpha h/2$. By Lemma˜6.4, we have $\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}\leq mL^{2}/\alpha$. Combining the above estimates finishes the proof.
5.6 Proof of Claim 5.1
Choose any $\varepsilon>0$. We have
$-2[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]\leq\varepsilon h|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+\frac{|A^{i,j}|^{2}}{\varepsilon h}$
and thus,
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{-1}\mathbb{E}|A^{i,j}|^{2}.$ (5.12)
By the definition of $A^{i,j}$ and Lemma˜6.4,
$|A^{i,\tau_{n,i}^{-1}(j)}|\leq\int_{nh}^{(n+1)h}|\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{nh}^{i,j},q_{*}^{-i})|\mathrm{d}t\leq\int_{nh}^{(n+1)h}L|Y_{t}^{i,j}-Y_{nh}^{i,j}|\mathrm{d}t.$
Then,
$\mathbb{E}|A^{i,\tau_{n,i}^{-1}(j)}|^{2}$ $\leq\mathbb{E}\bigg(\int_{0}^{h}L|Y_{nh+t}^{i,j}-Y_{nh}^{i,j}|\mathrm{d}t\bigg)^{2}=L^{2}h^{2}\mathbb{E}\bigg(\frac{1}{h}\int_{0}^{h}|Y_{nh+t}^{i,j}-Y_{nh}^{i,j}|\mathrm{d}t\bigg)^{2}$
$\leq L^{2}h^{2}\mathbb{E}\bigg(\frac{1}{h}\int_{0}^{h}|Y_{nh+t}^{i,j}-Y_{nh}^{i,j}|^{2}\mathrm{d}t\bigg)=L^{2}h\int_{0}^{h}\mathbb{E}|Y_{nh+t}^{i,j}-Y_{nh}^{i,j}|^{2}\mathrm{d}t.$ (5.13)
By (5.4), we have
$|Y_{nh+t}^{i,j}-Y_{nh}^{i,j}|^{2}$ $=\bigg(-\int_{0}^{t}\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})\mathrm{d}s+\sqrt{2}(B_{nh+t}^{i,\tau_{n,i}^{-1}(j)}-B_{nh}^{i,\tau_{n,i}^{-1}(j)})\bigg)^{2}$
$\leq 2\bigg[\bigg(\int_{0}^{t}\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})\mathrm{d}s\bigg)^{2}+[\sqrt{2}(B_{nh+t}^{i,\tau_{n,i}^{-1}(j)}-B_{nh}^{i,\tau_{n,i}^{-1}(j)})]^{2}\bigg].$ (5.14)
@@ 5.5.5 Step 5: Proving the recursive inequality
$\bigg(\int_{0}^{t}\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})\mathrm{d}s\bigg)^{2}=t^{2}\bigg(\frac{1}{t}\int_{0}^{t}\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})\mathrm{d}s\bigg)^{2}\leq t\int_{0}^{t}|\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})|^{2}\mathrm{d}s.$ (5.15)
Since $Y_{nh+s}^{:,j}\sim q_{*}$, we have
$\mathbb{E}|\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})|^{2}$ $=\mathbb{E}\Big|\mathbb{E}_{\bm{y}\sim q_{*}^{-i}}[\partial_{i}V(Y_{nh+s}^{i,j},\bm{y})]\Big|^{2}=\mathbb{E}\Big|\mathbb{E}[\partial_{i}V(\bm{Y}_{nh+s}^{:,j})|Y_{s}^{i,j}]\Big|^{2}$
$\leq\mathbb{E}|\partial_{i}V(\bm{Y}_{nh+s}^{:,j})|^{2}=\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}.$ (5.16)
By (5.14), (5.15) and (5.16),
$\mathbb{E}|Y_{nh+t}^{i,j}-Y_{nh}^{i,j}|^{2}$ $\leq 2t\int_{0}^{t}\mathbb{E}|\bar{V}_{i}^{\prime}(Y_{nh+s}^{i,j},q_{*}^{-i})|^{2}\mathrm{d}s+4\mathbb{E}|B_{nh+t}^{i,\tau_{n,i}^{-1}(j)}-B_{nh}^{i,\tau_{n,i}^{-1}(j)}|^{2}$
$\leq 2t^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+4t.$
Combining this with (5.13), we obtain that
$\mathbb{E}|A^{i,\tau_{n,i}^{-1}(j)}|^{2}\leq L^{2}h\int_{0}^{h}\Big(2t^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+4t\Big)\mathrm{d}t=\frac{2L^{2}h^{4}}{3}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+2L^{2}h^{3}.$
Plugging this into (5.12) yields the claimed result.